OpenStudy (anonymous):

Solve the following system by using graph paper or graphing technology. 2x + 2y = –6 3x – 2y = 11 What is the solution to the system?

OpenStudy (anonymous):

Each equation represents a line. Graph them both and the point where they cross gives the coordinates of the solution. The x,y coordinates are the x and y values that satisfy both equations.

OpenStudy (anonymous):

(–1, –7) (1, –4) (2, –1) (3, –2)

OpenStudy (anonymous):

If you don't need to show your work (the actual graph) just substitute each solution pair into the equations and whichever pair satisfies both is the solution.

OpenStudy (anonymous):

like i kinda of understand but i get it wrong

OpenStudy (anonymous):

2(-1)+2(-7)=-6 First option 2(1)+2(-4)=-6 Second 2(2)+2(-1)=-6 Third 2(3)+2(-2) = -6 Fourth Solve the left side on each option. Which one comes out as -6?

OpenStudy (anonymous):

Next, take the values you now think of the solution and substitute them also on the second equation. If they give the correct answer, you're done.